OpenStudy (anonymous):

solve the equation 2x+3=9

OpenStudy (anonymous):

Do you know what to do?

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

Frost you subtract 3 from each side and it becomes 2x. = 9 then you divide and simplify

OpenStudy (anonymous):

X=3

OpenStudy (anonymous):

No its 6

OpenStudy (anonymous):

2*6=12 12-3 =9 2*3=6 6-3=3 The answer is 6

OpenStudy (anonymous):

No because you subtract 3 from each side and the it's 2x=6 then you simplfy to. X=3

OpenStudy (anonymous):

It's 2x not two times

OpenStudy (anonymous):

U got it @Charlayah_Ridgell ?

OpenStudy (anonymous):

2x is 2 times x dumbass/ -,-

OpenStudy (anonymous):

But the answer is three. I thought the equation said -3 not +3

OpenStudy (anonymous):

yea 2 times x not two times 3 smart one